Frances McNale is an employee who drives a 2017 Ford C-Max with a fair-market value of $29,500. The lease value is $7,750, according to Publication 15-b. The car was available for 150 days during the year, and she reported that 6,500 of the 27,850 miles were driven for personal purposes. The company pays for all fuel and is charged back to Frances $0.055 per mile. What is the amount of the company-car fringe benefit that will appear on Frances's W-2, using the lease-value rule? 
Answer #1

Ans: Annual Value lease= 7,750

Days available for use=150

Total days in a year= 365

Prorated annual lease= 150/365*100

= 41.10%

Prorated annual lease= 41.10%* 7,750

= $3,184.93

Percentage for personal use during the year= 6500/27,850

=23.34%

Personal Annual lease value= 3184.93*23.34%

= $ 743.34

Fuel Value= 357.50 {6500*0.055}

Total Value that will appear= $743.34+357.50

= $ 1,100.84